OpenStudy (anonymous):

show that (the integral of x/(3x^2+3) from x=5 to x=7 actually equals (1/6)[ln(150)-ln(78)]...the closest I can get is to show that the integral is equal to (1/2)[ln(150)-ln(78)] and I'm on the verge of thinking my teacher is wrong, however when I put both of these into wolfram, they came out to be o.11108 and 0.10898, which is very close, so please anyone...HELP ME BEFORE I GO CRAZY!

OpenStudy (zehanz):

U can use the substitution u = 3x²+3. Then du = 6x. Now write the integral as follows:\[\int\limits_{5}^{7}\frac{ x }{ 3x^2+3 }dx=\frac{ 1 }{ 6 } \int\limits_{5}^{7}\frac{ 6x }{ 3x^2+3 }dx=\int\limits_{?}^{?}\frac{ 1 }{ u }du\] This one is much easier to find! Just one problem left: calculate the new boundaries for u. So: if x=5 then u=? if x=7 then u=?

OpenStudy (zehanz):

Ah, you probably did all this stuff already ;) So:\[\frac{ 1 }{ 6 }\int\limits_{78}^{150}\frac{ 1 }{ u }du=\frac{ 1 }{ 6 }\left( \ln 150 - \ln 78 \right)\], the same as your result!

OpenStudy (zehanz):

Was it the boundaries?

OpenStudy (anonymous):

hi...the computer is taking your formulas and making them very hard to read

OpenStudy (zehanz):

OK, I will upload it as an image...

OpenStudy (anonymous):

when i did it, i made u=x^2+1, which made 1/2du=x dx but i kept the "3" that i factored out of the deominator on the bottom...then i took the integral of ln(1/(3u) du

OpenStudy (anonymous):

then that made 1/2[3(x^2+1)...then i replaced x=50 and x=26

OpenStudy (anonymous):

i saw your image...i need a few minutes to look at it and understand it...

OpenStudy (anonymous):

So ZeHanz...were you actually able to get the left to equal the right?

OpenStudy (zehanz):

But if you factor out the 3, it can be put in front of the integral sign as 1/3. You then would have:\[\frac{ 1 }{ 3 }\int\limits_{5}^{7}\frac{ x }{ x^2+1 }dx=\frac{ 1 }{ 6 }\int\limits_{26}^{50}\frac{ 1 }{ u }du\] where u=x²+1 and du=2xdx, so the result would be: 1/6(ln(50)-ln(26))

OpenStudy (zehanz):

No matter what you do, you always get the same answer: 1/6(ln(150-ln(78))=1/6(ln(50)-ln(26))=1/6(ln(25)-ln(13)), because 150/78=50/26=25/13.

OpenStudy (anonymous):

hi ZeHanz...i really appreciate your help! Is there any way you could make an image of the entire problem? I can't understand what the computer has made of your entries

OpenStudy (zehanz):

OK, I'll do the same image-trick again!

OpenStudy (zehanz):

Here it is:

OpenStudy (anonymous):

oh wait a second...1/6[ln(150)-ln(78)] is equal to 1/6(ln(50)-ln(26))???...i think i had that second part as an answer but didn't know that it equaled what i was looking for

OpenStudy (anonymous):

yeah...i can get 1/6 out in front, but then i am removing the "3" that i need to multiply with both "50" and "26", and i then cannot get the 150 and 78 that i need

OpenStudy (anonymous):

would you do me a huge favor and show it all worked out for me? i still can't get it to work unfortunately...i have spent nearly 3 hours on this problem and i am going insane!...please help if you can

OpenStudy (anonymous):

Hey ZeHanz...I finally got it!!!!!!!!!!! THANK YOU THANK YOU THANK YOU!

OpenStudy (zehanz):

Here is my calculation...

OpenStudy (anonymous):

Hi ZeHanz, I just thought you would be interested to know that my teacher will only accept 1/6[ln(150)-ln(78)] as an answer...when I made u=x^2+1, i could only get the answer 1/6[ln(50)-ln(26)]...and even though that is mathematically equal to 1/6[ln(150)-ln(78)], he would have marked me wrong...my teacher is very tricky because the only way you can get 1/6[ln(150)-ln(78)], is to set u=3x^2+3...this is very tricky problem but I wanted to say thank you again...you saved my sanity!!! Danke!